OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 25 y = 2x − 5 −4 and 0 4 and 0 −3 and −5 3 and −5

OpenStudy (danjs):

it is a circle and a line, so it will intersect never, once as a tangent, or twice as a secant through th ecircle

OpenStudy (danjs):

They give you the value for y, y=2x-5 put that into the circle equation, and solve for x \[x^2+(2x-5)^2=25\]

OpenStudy (anonymous):

i don't know how to do that can you please help

OpenStudy (danjs):

\[x^2+(2x-5)(2x-5)=25\] \[x^2+4x^2-10x -10x+25=25\]

OpenStudy (anonymous):

I'm following

OpenStudy (danjs):

\[5x^2-20x = 0\] \[5x(x-4) = 0\] \[5x=0~~~~~or~~~~~x-4=0\]

OpenStudy (anonymous):

I'm following

OpenStudy (danjs):

if you are following till now , you can finish er

OpenStudy (anonymous):

so would it be -4 and 0?

OpenStudy (anonymous):

@DanJS